(a*1 1/2) +a=250 I can't figure out what to do with 1 1/2

Answer by edjones(8007)

(a * 1 1/2) + a = 250
3a/2 + a = 250 [1 1/2 = 3/2]
3a+2a=500 multiply each side by 2.
5a=500
a=100
.
Ed